A woman earns 15% more than her husband. Together they make $69300 per year. What is her monthly salary?
Answer by mananth(16949)

let husband earn % x
wife earns 1.15x
..
x+1.15x= 69300
2.15x=69300
x= 69300/2.15
x= 32,232 husband's income
wife = 1.15 * 32232
=37,068
...
wife annual salary /12 = 37,068/12
$ 3089 / month
...
m.ananth@hotmail.ca